estimate the terminal value of a company if the discount rate is 25%, annual growth rate for next year to infinity is 6% and the current years free cash flow is $612000

If the company had calculated the terminal value and the end of six years as $4100000, what would be its present value give same 25% discount rate?
Answer #1

(a) Discount Rate = 25 %, Annual Growth Rate = 6 %, Current Free Cash Flow (FCF0) = $ 612000

Expected FCF Next Year = FCF1 = 612000 x 1.06 = $ 648720

Terminal Value = 648720 / (0.25 - 0.06) = $ 3414315.789

(b) Terminal Value at the end of Year 6 = $ 4100000 and Discount Rate = 25 %

Present Value = 4100000 / (1.25)^(6) = $ 1074790.4